# GPU-based compressible lattice Boltzmann simulations on non-uniform grids using standard C++ parallelism: From best practices to aerodynamics, aeroacoustics and supersonic flow simulations

### v2

@@ Abstract
Despite decades of research, creating accurate, robust, and efficient lattice Boltzmann methods (LBM) on non-uniform grids with seamless GPU acceleration remains challenging. This work introduces a novel strategy to address this challenge by integrating simple yet effective components: (1) parallel algorithms in modern C++, (2) conservative cell-centered grid refinement, (3) local boundary conditions, and (4) robust collision models. Our framework supports multiple lattices (D2Q9, D2Q13, D2Q21, D2Q37) tailored to various flow conditions. It includes collision models with polynomial and numerical equilibria, a second distribution for polyatomic behavior, a Jameson-like shock sensor, and generalizes Rohde’s refinement strategy. The framework’s accuracy and robustness is validated across diverse benchmarks, including lid-driven cavity flows, Aeolian noise, 30P30N airfoil aerodynamics, inviscid Riemann problems, and viscous flows past a NACA airfoil in transonic and supersonic regimes. Modern C++ further enables our framework to reach GPU-native performance, while ensuring high portability, modularity, and ease of implementation. Notably, weakly compressible LBMs achieve state-of-the-art GPU efficiency on non-uniform grids, while fully compressible LBMs benefit from acceleration equivalent to thousands of CPU cores in the most compute-intensive cases. Our advanced performance models incorporate neighbor-list and asynchronous time-stepping effects, providing new insights into the performance decomposition of LB simulations on non-uniform grids. Overall, this study sets a new standard for portable, tree-based LBMs, demonstrating that a combination of well-chosen components can achieve high performance, accuracy, and robustness across various flow conditions. As a final proof-of-concept, adaptive mesh refinement is proposed for subsonic and supersonic applications.
@@ 1 Introduction
Fig 1: 2D illustration of the three families of grid refinement strategies available in the LB context [13, 14]. From left to right: vertex-centered (VC), cell-centered (CC), and combined (CM) approaches. Filled black circles and unfilled circles indicate locations where fine and coarse LB data is stored, respectively. When circles overlap, as in the vertex-centered approach, LB data is available at both coarse and fine levels.
Without delving too deeply into technical details, the reconstruction of missing information at the interface between two refinement levels requires tackling two primary challenges: (i) violation of exact conservation rules and (ii) improper rescaling of LB data. Most grid refinement algorithms based on VC and CM approaches pay much attention to the second issue, often overlooking accuracy problems caused by improper mass, momentum, and energy conservation at the interface. Conversely, CC strategies, such as the one proposed by Rhode et al. [24], automatically enforce conservation rules. The conservation issue was only recently investigated by Astoul et al. [21, 2] for VC algorithms, and more broadly by Schukmann et al. [13, 14] for CC, VC, and CM algorithms. Aggregating the results of these studies, it appears that ensuring exact conservation rules at the interface significantly improves accuracy. This property brings back Rhode’s approach [24] to the forefront of discussions, as, additionally to its natural conservation properties, the implementation of the model is particularly straightforward and does not require any explicit interpolation scheme, or handling of special cases near boundaries. Furthermore, it can easily be extended to any velocity discretization, which is essential for simulating supersonic flows with high-order LBMs.. All of this leads to the presumption, further substantiated by the results presented in this article, that Rhode’s approach and its high-order extension are very good candidates for the efficient implementation of general LB solvers on non-uniform grids.
@@
The paper is organized as follows. Sections 2 and 3 provide a brief overview of LBMs and a detailed explanation of our grid-refinement strategy. Section 4 discusses implementation details, focusing on the use of parallel algorithms for the portable execution of tree-based LBMs on both CPUs and GPUs. Validation of the framework is presented in Sections 5 and 6, using well-established benchmarks for low-speed and high-speed flows in 2D, respectively. Performance analyses are conducted in Section 7 for both weakly and fully compressible LBMs. Section 8 presents preliminary results on adaptive mesh refinement for subsonic and supersonic flow simulations using high-order LBMs. Conclusions and perspectives are addressed in Section 9.
@@ 2 General background on lattice Boltzmann methods
The latter describes the space and time evolution of $V$ groups of fictive particles through their distribution function $f_{i}$ ($i\in\{0,...,V-1\}$), which can be interpreted as the number of fictive particles located at $(\bm{x},t)$ and propagating at a mesoscopic velocity $\bm{\xi}_{i}$. These particles are assumed to either collide or stream on a Cartesian grid, leading to the efficient collide-and-stream algorithm [58]
@@ 3.2 Present grid refinement framework
Fig 3: Illustration of the buffer layers (in grey) required by CC approaches for standard and high-order lattices. Circles and squares correspond to fine and coarse nodes respectively. Red nodes are used to transfer data between coarse and fine levels through the coalescence and explosion steps [24].
@@ 4.4 Data layout for efficient GPU acceleration on trees
The final aspect requiring attention is the method for accessing neighbor data during the streaming step. One approach consists of traversing the tree data structure for every neighbor access, while storing a single tree-node index per cell as an entry point for tree traversal. However, preliminary tests revealed that the computational cost of repeatedly executing a tree-search algorithm remains prohibitively high on modern GPUs, despite their growing support for complex execution flows. Instead, we propose to compute the neighbor list during the preprocessing step. This involves generating the Cartesian coordinates of all neighboring cells and searching the tree for cells that match these coordinates. The SFC index of the identified cells is in a direct relationship with the storage index of the neighboring cell’s data, which is subsequently stored as a component of the neighbor list. Thus, excluding the rest population, $V-1$ integer values are stored for every cell along with the $V$ populations. We store these neighbor indices as 32-bit integers, hence allowing to handle meshes with up to 4 billion cells, which is sufficient for GPUs with current memory capability. Given the continuous increase of performance and memory size of GPUs, it must however be expected to switch to 64-bit integers in a near future, which will automatically double the cost of these memory accesses. The neighbor lists are eventually stored in a AoS format as they are accessed locally on every cell.
@@
    Each cell also stores $V-1$ neighbor links as 32-bit integers in an AoS layout. The AoS format is chosen because neighbor lists are accessed locally for each cell during the streaming step.
@@ 5.2 Steady internal aerodynamics: Lid-driven cavity
To assess convergence of the simulation, the total kinetic energy in the domain, $kin$, is monitored. When the normalized difference $|kin(t_{1})-kin(t_{0})|/kin(t_{1})$ reaches $10^{-6}$, the flow is considered to have reached a steady state. This global criterion is evaluated each $t_{1}-t_{0}=0.1t_{c}$, with the convective time being $t_{c}=L/u_{\infty}$. Figure 7 shows the local Mach fields obtained after convergence, for each mesh configuration, using the D2Q9-RR model, and for $\mathrm{Re}=100$. Qualitatively, all configurations result in a primary recirculation of similar size and location, and this is also true for high-order lattices (not shown here). No accuracy issues are observed near refinement interfaces or at flat/corner boundary conditions. To verify the accuracy of our refinement strategy, we compare velocity profiles along the centerlines with reference data from Ghia et al.[116]. Remarkably, all lattices and mesh configurations show excellent agreement with the reference data (see Figure 8). In particular, the vertical velocity profiles measured within the buffer layers for configuration "vert" confirm the accurate reconstruction of missing information, even when using two (D2Q13) or three (D2Q21) consecutive buffer layers.
@@ 5.3 Aeolian aeroacoustics: Flow past a cylinder
Fig 10: Aeolian noise characteristics at $(\mathrm{Re},\mathrm{Ma},\Delta x_{f})=(150,0.2,D/160)$ using the D2Q9-RR-LBM. Left panel: Time evolution of pressure fluctuations, $p^{\prime}=(p-p_{\mathrm{mean}})/p_{\infty}$, at $r/D=75$ and for $\theta\in\{+90^{\mathrm{o}},180^{\mathrm{o}},-90^{\mathrm{o}}\}$. Middle panel: Time evolution of normalized pressure fluctuations, $p^{\prime}_{\mathrm{norm}}=p^{\prime}/p_{\infty}(r/r_{\mathrm{ref}})^{1/2}$, for various locations $r/D\in\{50,75,100\}$ and $\theta=\pm 90^{\mathrm{o}}$. Right panel: Directivity of the normalized root-mean-square (rms) pressure, $p^{\mathrm{rms}}_{\mathrm{norm}}=p^{\mathrm{rms}}/p_{\infty}(r/r_{\mathrm{ref}})^{1/2}$. The reference distance, $r_{\mathrm{ref}}/D=75$, is chosen to facilitate comparison with DNS data presented in Figure 12 of Ref. [119]. Mean and rms values are calculated over the final $100t_{c}$. $t_{s}$ accounts for the time shift between different locations, assuming sound waves propagate at the speed of sound $c_{s}$.
@@
Figure 10 gathers more quantitative data about the dipolar nature of the sound radiated from the cylinder. For such configuration, lift and drag fluctuations are expected to lead to pressure waves of different amplitudes and frequency [119, 120]. These fluctuations are captured by the pressure measurements at $r/D=75$ for various angles $\theta\in\{+90^{\mathrm{o}},180^{\mathrm{o}},-90^{\mathrm{o}}\}$. More precisely, lift force fluctuations in the ±$\pm 90^{\mathrm{o}}$ directions produce acoustic waves with amplitudes approximately 10 times larger than those generated by drag fluctuations in the $180^{\mathrm{o}}$ direction. In our simulations, the lift dipole oscillates at half the frequency of the drag dipole. The Strouhal number was determined via FFT analysis of the time signal recorded over $100t_{c}$, utilizing Hamming windowing and zero-padding to enhance resolution [127]. This method yielded $\mathrm{St}=0.1824$, which aligns closely with DNS results [119] and experimental data [128]: $\mathrm{St}=0.183$ and $\mathrm{St}=0.179-0.185$, respectively. These findings remain consistent across different observation points ($r/D=50$, $r/D=75$, and $r/D=100$), as illustrated in the middle panel of Figure 10. Interestingly, the pressure fluctuations exhibit a form of “self-similar” behavior as all curves almost perfectly coincide when (1) their amplitude is normalized by the decay rate of $1/r^{2}$, and (2) their time evolution is shifted according to the speed of sound and the distance between the two locations. The right panel of Figure 10 further quantifies the noise directivity using the normalized root-mean-square (rms) pressure values, $p^{\mathrm{rms}}_{\mathrm{norm}}$. The observed rms values, particularly the maxima in the ±$\pm 90^{\mathrm{o}}$ directions, are in excellent agreement with DNS data from Inoue et al. [119]. This polar plot confirms that, regardless of the measurement distance, the emitted noise is predominantly dipolar in nature, with the lift dipole serving as the primary contributor to the radiated acoustic field. In fact, this holds true for a wide range of Reynolds and Mach numbers as shown in B.3.
@@ 5.4 External aerodynamics at realistic conditions: Flow past the 30P30N three-element airfoil
To investigate the ability of our framework in simulating high-lift configurations under realistic conditions, we conducted simulations of the flow past a 30P30N three-element airfoil at $(\mathrm{Re},\mathrm{Ma})=(1.7\times 10^{6},0.17)$ using the D2Q9-RR-LBM. For angles of attack (AoAs) considered in this work, $\mathrm{AoA}\in\{5.5^{\circ},9.5^{\circ},14^{\circ},20^{\circ}\}$, the local Mach number reaches values of 0.5-0.6 in between the slat and main wing, and this pushes the D2Q9-LBM to its stability limit. For the highest AoA, the boundary layer becomes increasingly sensitive to no-slip boundary conditions, often leading to early detachment in case of inaccurate wall treatment. This provides valuable insights into scenarios where stair-cased boundaries, such as bounce-back conditions, may require improvements (e.g., interpolation or wall modeling) to capture the correct physics. In the end, these extreme conditions are a formidable opportunity to assess the stability and accuracy limits of choices made in this work, to provide best practices to the community, and and to guide future research.
@@
To further improve the stability of the D2Q9-RR-LBM, we chose to equilibrate the trace of the second-order moments, a method widely used in multi-relaxation-time and central-moment-based collision models for single-phase [143, 144, 145], multiphase [146, 147], and magnetohydrodynamic flow simulations [148, 149]. This introduces an additional bulk viscosity proportional to $\Delta t$ [150]. As such it adapts to the local mesh resolution, hence being smaller in well-resolved regions and larger in more critical under-resolved areas. The method is generalizable to any collision model [64], with implementations available in open-source solvers like Palabos [97, 12] and STLBM [11].
@@ 6.2 Inviscid 2D Riemann problem
To further investigate the ability of our refinement strategy to accurately and robustly simulate the above phenomena in inviscid conditions ($\nu=\nu_{T}=0$), we consider three mesh configurations to discretize the $[-L,L]\times[-L,L]$ simulation domain: diag, axes, and optim. The diag configuration features a grid refinement interface along the diagonal $x=y$, hence, aligning with the symmetry plane of the initial state. This setup was chosen because it poses significant challenges in terms of accuracy and robustness, as various physical phenomena evolve along this diagonal. The configuration axes employs refinement patches along the mesh axes, which enhances the capture of the initial evolution of discontinuities. However, this configuration is less effective at accurately capturing the interaction and evolution of waves along the diagonal axis ($x=y$). Lastly, optim was designed to represent an optimal configuration that successfully captures the initial evolution of discontinuities, the complex patterns emerging in quadrant [1], as well as the mushroom-like vortices forming in quadrant [3].
Neumann boundary conditions are unsuitable here because the diag configuration has a refinement interface that intersects the boundaries in a non-axis-aligned manner, which the current version of the code cannot perfectly handle. Instead, periodic boundary conditions are applied at all boundaries and data is outputted on a reduced domain $[-L/2,L/2]\times[-L/2,L/2]$ at $t=0.30t_{c}$, where the characteristic time is $t_{c}=L/\sqrt{\gamma_{r}T_{0}}$. On top of the three mesh configurations, three mesh resolutions are considered to better quantify the accuracy and robustness of our approach in severely under-resolved ($N_{c}=32$) and well-resolved ($N_{c}=512$) conditions.

### v3

@@ Abstract
Despite decades of research, creating accurate, robust, and efficient lattice Boltzmann methods (LBM) on non-uniform grids with seamless GPU acceleration remains challenging. This work introduces a novel strategy to address this challenge by integrating simple yet effective components: (1) parallel algorithms in modern C++, (2) conservative cell-centered grid refinement, (3) local boundary conditions, and (4) robust collision models. Our framework supports multiple lattices (D2Q9, D2Q13, D2Q21, D2Q37, D3Q27, etc) tailored to various flow conditions. It includes collision models with polynomial and numerical equilibria, a second distribution for polyatomic behavior, a Jameson-like shock sensor, and generalizes Rohde’s refinement strategy.
The framework’s accuracy and robustness is validated across diverse benchmarks, including lid-driven cavity flows, Aeolian noise, 30P30N airfoil aerodynamics, inviscid Riemann problems, and viscous flows past a NACA airfoil in transonic and supersonic regimes. Modern C++ further enables our framework to reach GPU-native performance, while ensuring high portability, modularity, and ease of implementation. Notably, weakly compressible LBMs achieve state-of-the-art GPU efficiency on non-uniform grids, while fully compressible LBMs benefit from acceleration equivalent to thousands of CPU cores in the most compute-intensive cases. Our advanced performance models incorporate neighbor-list and asynchronous time-stepping effects, providing new insights into the performance decomposition of LB simulations on non-uniform grids.
Overall, this study sets a new standard for portable, tree-based LBMs, demonstrating that a combination of well-chosen components can achieve high performance, accuracy, and robustness across various flow conditions. As a final proof-of-concept, adaptive mesh refinement is proposed for subsonic and supersonic applications.
@@ 1 Introduction
Fig 1: 2D illustration of the three families of grid refinement strategies available in the LB context [13, 14]. From left to right: vertex-centered (VC), cell-centered ($C$), and combined (CM) approaches. Filled black circles and unfilled circles indicate locations where fine and coarse LB data is stored, respectively. When circles overlap, as in the vertex-centered approach, LB data is available at both coarse and fine levels.
Without delving too deeply into technical details, the reconstruction of missing information at the interface between two refinement levels requires tackling two primary challenges: (i) violation of exact conservation rules and (ii) improper rescaling of LB data. Most grid refinement algorithms based on VC and CM approaches pay much attention to the second issue, often overlooking accuracy problems caused by improper mass, momentum, and energy conservation at the interface. Conversely, CC strategies, such as the one proposed by Rohde et al. [24], automatically enforce conservation rules. The conservation issue was only recently investigated by Astoul et al. [21, 2] for VC algorithms, and more broadly by Schukmann et al. [13, 14] for CC, VC, and CM algorithms. Aggregating the results of these studies, it appears that ensuring exact conservation rules at the interface significantly improves accuracy. This property brings back Rohde’s approach [24] to the forefront of discussions, as, additionally to its natural conservation properties, the implementation of the model is particularly straightforward and does not require any explicit interpolation scheme, or handling of special cases near boundaries. Furthermore, it can easily be extended to any velocity discretization, which is essential for simulating supersonic flows with high-order LBMs.. All of this leads to the presumption, further substantiated by the results presented in this article, that Rohde’s approach and its high-order extension are very good candidates for the efficient implementation of general LB solvers on non-uniform grids.
@@
The paper is organized as follows. Sections 2 and 3 provide a brief overview of LBMs and a detailed explanation of our grid-refinement strategy. Section 4 discusses implementation details, focusing on the use of parallel algorithms for the portable execution of tree-based LBMs on both CPUs and GPUs. Validation of the framework is presented in Sections 5 and 6, using well-established benchmarks for low-speed and high-speed flows in 2D, respectively. Performance analyses are conducted in Section 7 for both weakly and fully compressible LBMs. Section 8 presents preliminary results on adaptive mesh refinement for subsonic and supersonic flow simulations using high-order LBMs. A summary of our findings is proposed in Section 9. Future research directions are addressed in Section 10 and illustrated with the 3D simulation of the turbulent flow past a sphere.
@@ 2 General background on lattice Boltzmann methods
The latter describes the space and time evolution of $V$ groups of fictive particles through their distribution function $f_{i}$ ($i\in\{0,...,V-1\}$), which can be interpreted as the mass density of fictive particles located at $(\bm{x},t)$ and propagating at a mesoscopic velocity $\bm{\xi}_{i}$. These particles are assumed to either collide or stream on a Cartesian grid, leading to the efficient collide-and-stream algorithm [58]
@@ 3.2 Present grid refinement framework
Fig 3: Illustration of the buffer layers (in grey) required by $C$ approaches for standard and high-order lattices. Circles and squares correspond to fine and coarse nodes respectively. Red nodes are used to transfer data between coarse and fine levels through the coalescence and explosion steps [24].
@@ 4.4 Data layout for efficient GPU acceleration on trees
The final aspect requiring attention is the method for accessing neighbor data during the streaming step. One approach consists of traversing the tree data structure for every neighbor access, while storing a single tree-node index per cell as an entry point for tree traversal. However, preliminary tests revealed that the computational cost of repeatedly executing a tree-search algorithm remains prohibitively high on modern GPUs, despite their growing support for complex execution flows. Instead, we propose to compute the neighbor list during the preprocessing step. This involves generating the Cartesian coordinates of all neighboring cells and searching the tree for cells that match these coordinates. The SFC index of the identified cells is in a direct relationship with the storage index of the neighboring cell’s data, which is subsequently stored as a component of the neighbor list. Thus, excluding the rest population, V1V-1 integer values are stored for every cell along with the $V$ populations. We store these neighbor indices as 32-bit integers, hence allowing to handle meshes with up to 4 billion cells, which is sufficient for GPUs with current memory capability. Given the continuous increase of performance and memory size of GPUs, it must however be expected to switch to 64-bit integers in a near future, which will automatically double the cost of these memory accesses. The neighbor lists are eventually stored in a AoS format as they are accessed locally on every cell.
@@
    Each cell also stores V1V-1 neighbor links as 32-bit integers in an AoS layout. The AoS format is chosen because neighbor lists are accessed locally for each cell during the streaming step.
@@ 5.2 Steady internal aerodynamics: Lid-driven cavity
To assess convergence of the simulation, the total kinetic energy in the domain, kinkin, is monitored. When the normalized difference $|kin(t_{1})-kin(t_{0})|/kin(t_{1})$ reaches $10^{-6}$, the flow is considered to have reached a steady state. This global criterion is evaluated each $t_{1}-t_{0}=0.1t_{c}$, with the convective time being $t_{c}=L/u_{\infty}$. Figure 7 shows the local Mach fields obtained after convergence, for each mesh configuration, using the D2Q9-RR model, and for $\mathrm{Re}=100$. Qualitatively, all configurations result in a primary recirculation of similar size and location, and this is also true for high-order lattices (not shown here). No accuracy issues are observed near refinement interfaces or at flat/corner boundary conditions. To verify the accuracy of our refinement strategy, we compare velocity profiles along the centerlines with reference data from Ghia et al.[116]. Remarkably, all lattices and mesh configurations show excellent agreement with the reference data (see Figure 8). In particular, the vertical velocity profiles measured within the buffer layers for configuration vert confirm the accurate reconstruction of missing information, even when using two (D2Q13) or three (D2Q21) consecutive buffer layers.
@@ 5.3 Aeolian aeroacoustics: Flow past a cylinder
Fig 10: Aeolian noise characteristics at $(\mathrm{Re},\mathrm{Ma},\Delta x_{f})=(150,0.2,D/160)$ using the D2Q9-RR-LBM. Left panel: Time evolution of pressure fluctuations, $p^{\prime}=(p-p_{\mathrm{mean}})/p_{\infty}$, at $r/D=75$ and for $\theta\in\{+90^{\mathrm{o}},180^{\mathrm{o}},-90^{\mathrm{o}}\}$. Middle panel: Time evolution of normalized pressure fluctuations, $p^{\prime}_{\mathrm{norm}}=p^{\prime}/p_{\infty}(r/r_{\mathrm{ref}})^{1/2}$, for various locations $r/D\in\{50,75,100\}$ and $\theta=\pm 90^{\mathrm{o}}$. Right panel: Directivity of the normalized root-mean-square (rms) pressure, $p^{\mathrm{rms}}_{\mathrm{norm}}=p^{\mathrm{rms}}/p_{\infty}(r/r_{\mathrm{ref}})^{1/2}$. The reference distance, $r_{\mathrm{ref}}/D=75$, is chosen to facilitate comparison with DNS data presented in Figure 12 of Ref. [119]. Mean and rms values are calculated over the final $100t_{c}$. $t_{s}$ accounts for the time shift between different locations, assuming sound waves propagate at the speed of sound $c_{s}$. Reference $p^{\mathrm{rms}}$ values are extracted from Figure 12 (a) of Ref. [119] using WebPlotDigitizer.
@@
Figure 10 gathers more quantitative data about the dipolar nature of the sound radiated from the cylinder. For such configuration, lift and drag fluctuations are expected to lead to pressure waves of different amplitudes and frequency [119, 120]. These fluctuations are captured by the pressure measurements at $r/D=75$ for various angles $\theta\in\{+90^{\mathrm{o}},180^{\mathrm{o}},-90^{\mathrm{o}}\}$. More precisely, lift force fluctuations in the $\pm 90^{\mathrm{o}}$ directions produce acoustic waves with amplitudes approximately 10 times larger than those generated by drag fluctuations in the $180^{\mathrm{o}}$ direction. In our simulations, the lift dipole oscillates at half the frequency of the drag dipole. The Strouhal number was determined via FFT analysis of the time signal recorded over $100t_{c}$, utilizing Hamming windowing and zero-padding to enhance resolution [127]. This method yielded $\mathrm{St}=0.1824$, which aligns closely with DNS results [119] and experimental data [128]: $\mathrm{St}=0.183$ and $\mathrm{St}=0.179-0.185$, respectively. These findings remain consistent across different observation points ($r/D=50$, $r/D=75$, and $r/D=100$), as illustrated in the middle panel of Figure 10. Interestingly, the pressure fluctuations exhibit a form of “self-similar” behavior as all curves almost perfectly coincide when (1) their amplitude is normalized by the decay rate of $1/r^{2}$, and (2) their time evolution is shifted according to the speed of sound and the distance between the two locations. The right panel of Figure 10 further quantifies the noise directivity using the normalized root-mean-square (rms) pressure values, $p^{\mathrm{rms}}_{\mathrm{norm}}$. The observed rms values are in very good agreement with DNS data from Inoue et al. [119]. This polar plot confirms that, regardless of the measurement distance, the emitted noise is predominantly dipolar in nature, with the lift dipole serving as the primary contributor to the radiated acoustic field. In fact, this holds true for a wide range of Reynolds and Mach numbers as shown in B.3.
@@ 5.4 External aerodynamics at realistic conditions: Flow past the 30P30N three-element airfoil
To investigate the ability of our framework in simulating high-lift configurations under realistic conditions, we conducted simulations of the flow past a 30P30N three-element airfoil at $(\mathrm{Re},\mathrm{Ma})=(1.7\times 10^{6},0.17)$ using the D2Q9-RR-LBM. For angles of attack (AoAs) considered in this work, $\mathrm{AoA}\in\{5.5^{\circ},9.5^{\circ},14^{\circ},20^{\circ}\}$, the local Mach number reaches values of 0.5-0.6 in between the slat and main wing, and this pushes the D2Q9-LBM to its stability limit. For the highest AoA, the boundary layer becomes increasingly sensitive to no-slip boundary conditions, often leading to early detachment in case of inaccurate wall treatment. This provides valuable insights into scenarios where stair-cased boundaries, such as bounce-back conditions, may require improvements (e.g., interpolation or wall modeling) to capture the correct physics. In the end, these extreme conditions are a formidable opportunity to assess the stability and accuracy limits of choices made in this work, to provide best practices to the community, and to guide future research.
@@
To further improve the stability of the D2Q9-RR-LBM, we chose to equilibrate the trace of the second-order moments –a method similar to the bulk viscosity modification proposed by Dellar [108] which is widely used in multi-relaxation-time and central-moment-based collision models for single-phase [143, 144, 145], multiphase [146, 147], and magnetohydrodynamic flow simulations [148, 149]. This introduces an additional bulk viscosity proportional to $\Delta t$ [150]. As such it adapts to the local mesh resolution, hence being smaller in well-resolved regions and larger in more critical under-resolved areas. The method is generalizable to any collision model [64], with implementations available in open-source solvers like Palabos [97, 208] and STLBM [11].
@@ 6.2 Inviscid 2D Riemann problem
To further investigate the ability of our refinement strategy to accurately and robustly simulate the above phenomena in inviscid conditions ($\nu=\nu_{T}=0$), we consider three mesh configurations to discretize the $L,L]\times[-L,L]$ simulation domain: diag, axes, and optim. The diag configuration features a grid refinement interface along the diagonal $x=y$, hence, aligning with the symmetry plane of the initial state. This setup was chosen because it poses significant challenges in terms of accuracy and robustness, as various physical phenomena evolve along this diagonal. The configuration axes employs refinement patches along the mesh axes, which enhances the capture of the initial evolution of discontinuities. However, this configuration is less effective at accurately capturing the interaction and evolution of waves along the diagonal axis ($x=y$). Lastly, optim was designed to represent an optimal configuration that successfully captures the initial evolution of discontinuities, the complex patterns emerging in quadrant [1], as well as the mushroom-like vortices forming in quadrant [3].
Neumann boundary conditions are unsuitable here because the diag configuration has a refinement interface that intersects the boundaries in a non-axis-aligned manner, which the current version of the code cannot perfectly handle. Instead, periodic boundary conditions are applied at all boundaries and data is outputted on a reduced domain $L/2,L/2]\times[-L/2,L/2]$ at $t=0.30t_{c}$, where the characteristic time is $t_{c}=L/\sqrt{\gamma_{r}T_{0}}$. On top of the three mesh configurations, three mesh resolutions are considered to better quantify the accuracy and robustness of our approach in severely under-resolved ($N_{c}=32$) and well-resolved ($N_{c}=512$) conditions.